### v1

@@ 1 Introduction
Quantum computing and machine intelligence are converging to redefine how data-driven systems learn, adapt, and comply with privacy laws. As quantum machine learning (QML) transitions from small-scale prototypes to distributed deployments, the notion of the right to erasure acquires a new dimension. Under frameworks such as the GDPR, erasure must extend beyond model parameters to include quantum states, channels, and measurements [1, 2]. In this setting, unlearning is not a simple rollback, it is a physically valid transformation that decouples private signals from global quantum behaviour. Quantum mechanics offers natural privacy safeguards through no-cloning, measurement disturbance, and decoherence, yet these same principles make verification, audit, and decentralized enforcement challenging at scale.
@@
Federated learning (FL) decentralizes training but still exposes gradients and aggregates to inversion risks. Quantum Federated Learning (QFL) pushes privacy deeper into the computational stack by encoding updates as quantum states and coordinating learning through variational circuits [3]. When combined with quantum differential privacy (QDP), QFL can achieve rigorous confidentiality with minimal utility loss. Rofougaran et al. [4] demonstrate a QFL–QDP pipeline achieving $98\%$ accuracy under $\epsilon<1.3$ on NISQ hardware, showing that privacy and performance can coexist. This privacy posture now extends to the transport layer. Quantum cloud infrastructures and 6G space–ground networks rely on quantum key distribution (QKD) and homomorphic encryption to guarantee that deleted states remain computationally and physically unrecoverable. Zeng et al. introduce quantum circuit random reconstruction homomorphic encryption (QCRRA), enabling encrypted circuit execution even on untrusted servers [5]. Similarly, hybrid QFL combined with QKD improves both anomaly detection and key generation rates, indicating that privacy-preserving protocols can enhance, not hinder, performance [6].
@@
Quantum deployments also expand the threat model. Adversaries may conduct membership inference, inversion, or poisoning attacks on parametrized quantum circuits, exploiting universal perturbations that transfer across architectures [7]. A defensible system must reason about privacy at the level of quantum channels. A natural analogue of $(\varepsilon,\delta)$-differential privacy requires that, for neighbouring datasets encoded as states and for any POVM, the resulting output distributions differ by at most $(\varepsilon,\delta)$ after the full CPTP pipeline. Composition across rounds and clients then follows from channel algebra. In practice, calibrated dephasing or depolarizing noise acts as privacy randomization, while error mitigation restores performance. In federated setups, cryptographic transport is indispensable. Measurement-device-independent QKD provides composable keys for authenticated encryption and Pauli one-time pads during aggregation, as illustrated in smart-grid benchmarks [8]. Privacy must also persist through time. Quantum continual learning reduces catastrophic forgetting and constrains leakage from past tasks by allocating disjoint PQC subspaces or projecting gradients to avoid interference [9, 10]. Robustness-aware training, randomized encodings, and adversarial PQC optimization further limit attack transferability and restrict what can be inferred from exchanged statistics.
@@
Contributions. (1) A quantum information formalization of unlearning that replaces destructive erasure with CPTP-driven redistribution, anchored in the data-processing inequality and thermodynamic irreversibility. (2) A taxonomy and evaluation framework covering sample-, class-, and client-level scopes; empirical, certified, and DP guarantees; and mechanisms including reset, Fisher-weighted updates, and gradient reversal across QML and QFU deployments. (3) Federated QMU protocols that combine gradient hiding, differential privacy accounting, and cryptographic transport for secure orchestration [11, 12, 13, 14]. (4) Threat-informed design principles addressing transfer attacks and catastrophic forgetting through continual-learning and robustness-aware PQC training [7, 9, 10]. (5) Empirical evidence that privacy and utility can coexist in practice, including hybrid recurrent and time-series models where variational quantum embeddings improve forecasting with compact architectures [15].
@@ 2.1 From Classical to Quantum Unlearning
Machine unlearning (MU) aims to remove the influence of selected training data without retraining the entire model. The idea aligns with legal and ethical frameworks such as the European Union’s Right to Erasure in the GDPR [1, 2], which demand that data and its effects be verifiably erased. Early approaches used practical engineering strategies, including checkpoint rollback and incremental updates. Influence functions introduced a more principled approach by estimating each sample’s contribution to the optimum through $\mathbf{H}^{-1}\nabla_{\theta}L(\theta^{*},x_{i})$, with $\mathbf{H}$ as the empirical Hessian [16]. The SISA framework (Sharded, Isolated, Sliced Aggregation) further improved auditability and deletion latency through modular retraining [2].
Three common goals unify this body of work. Efficiency demands minimal computational cost and wall-clock time. Completeness ensures both data and its downstream influence are removed. Verifiability requires clear, measurable evidence of forgetting. Certified low-rank corrections enable exact updates in convex models [17], while approximate regularization and Bayesian re-weighting extend unlearning to deep, non-convex models by transitioning from $p(\theta|D)$ to $p(\theta|D\setminus D_{r})$ [18].
Federated unlearning applies similar ideas to distributed learning. Each client locally compensates for removed data by adjusting its gradient contribution, e.g., $\theta^{\prime}=\theta-\alpha\nabla_{\theta}L_{i}$, often using stored or compressed updates [19]. Hybrid quantum–classical systems now adopt comparable coordination strategies. Quantum feature extractors integrate with classical aggregators [20], ring topologies support bandwidth-efficient communication [21], and dynamic aggregation improves robustness under non-IID healthcare data [22]. Information-theoretically, ideal unlearning reduces mutual information between the removed data and the model, $I(D_{r};M)$, while retaining $I(D_{s};M)$ for the preserved set [23]. Metrics such as forgetting accuracy, retention, and retraining cost quantify progress toward this goal.
@@ 2.2 Quantum Learning Dynamics
which allows exact differentiation under real hardware noise [24].
Model expressivity depends on data encoding and circuit depth. Re-uploading strategies expand hypothesis space by alternating data injection with trainable layers [25], while differentiable quantum programming provides compiler-level support for gradient computation across control structures [26]. Training remains challenging due to barren plateaus, where gradient variance decays exponentially with circuit depth [27]. Geometry-aware optimization addresses this by using the quantum Fisher information (QFI) to rescale gradients, $\theta\leftarrow\theta-\eta F(\theta)^{-1}\nabla_{\theta}\mathcal{L}$, thereby improving convergence and shot efficiency on NISQ devices [28].
Distributed learning frameworks extend these foundations. In quantum federated learning (QFL), multiple clients train local PQCs or hybrid models and share updates under privacy-preserving protocols such as gradient hiding, quantum differential privacy, or entanglement-based coordination [11, 12, 13]. Comprehensive surveys document the architectural and privacy challenges that shape scalable QFL systems [14]. Collectively, exact gradient evaluation, QFI-guided optimization, and privacy-preserving coordination provide the algorithmic substrate on which QMU is built. They allow a quantum model to not only learn from data but also to forget it through controlled, measurable influence reduction.
@@ 2.3 Information Constraints: No-Cloning and No-Deletion
Logical erasure also has energetic cost. Landauer’s principle states that each bit of erased information dissipates heat [34, 35]. In QMU, forgetting increases entropy in the environment rather than annihilating stored information. For any completely positive trace-preserving (CPTP) map $\mathcal{E}$ and states $\rho,\sigma$, the data-processing inequality,
@@
ensures that distinguishability monotonically decreases under physical evolution [36, 37]. Trace distance and fidelity exhibit the same contractive behavior, forming the mathematical foundation for certified quantum forgetting.
@@
These physical limits redefine the purpose of unlearning. QMU cannot erase information; it instead makes the information operationally inaccessible. Practical pipelines combine three elements. First, CPTP channels provably contract trace distance or increase fidelity relative to the retrained counterfactual. Second, QFI-guided optimization reduces parameter sensitivity to individual samples. Third, federated and secure protocols prevent reconstruction through gradient hiding or differential privacy [38, 39, 40, 11, 41]. Adversarial analyses confirm that measurable contraction, not ideal erasure, defines robust unlearning under realistic noise [42, 43].
@@ 3.1 Scope: Sample, Class, and Client Levels
Scope specifies the unit of forgetting. At the sample level, the target set $\mathcal{S}=\{x_{j}\}$ contains individual records whose influence should be removed while preserving utility on retained data. In kernel QML, this often reduces to low-rank modifications of the Gram matrix $K_{jk}=\kappa_{\phi}(x_{j},x_{k})$ and efficient Sherman–Morrison–Woodbury updates; shallow, aligned kernels tend to dominate generalization and admit fast row/column re-estimation [44]. In variational models, edits localize to layers of $U(\theta)=\prod_{\ell}U_{\ell}(\theta_{\ell})$ guided by sensitivity; a first-order approximation
@@
captures the local effect of $\mathcal{S}$, with $H^{-1}$ replaced in practice by damped or block-diagonal QFI preconditioners for stability on NISQ hardware [45].
At the class level, we remove geometry associated with one or more labels and retain the rest. Mixed-state generative classifiers enable component-level deletion without retraining entire models [46]. Search over circuit families with tunable expressivity supports targeted suppression of class-specific features [47]. Compact QVC ensembles further allow pruning of subcircuits and light retraining of combiner layers for sequential tasks [48].
At the client level, relevant to QFU, forgetting combines parameter edits with state-theoretic maps. A local CPTP channel acts on the client register, followed by partial retraining of the global model; distributed QLSTM and kernelized pipelines implement such workflows with secure coordination [49, 50, 51]. Applications in security and materials show that lightweight hybrids with small latent spaces permit rapid reinitialization and refitting [52, 53, 54]. Choosing scope is thus a design decision: sample-level for precise removal [44], class-level for generative or structural shifts [46], and client-level when privacy or provenance dominates [49, 51]. Auditability and secure aggregation constraints should inform this choice [55, 56].
@@ 3.2 Guarantees: Empirical, Certified, and Differential Privacy
Guarantees make forgetting credible. Empirical guarantees rely on reproducible trends across seeds, backends, and datasets. QSVM and kernel hybrids show stable accuracy on compressed medical datasets such as CompressedMediQ and QProteoML [57, 58], and robustness persists under decoherence and backend variability [59, 60]. These studies validate observed forgetting behavior, although they do not bound worst cases.
@@
strengthened by precompression [57]. For variational models,
@@
and Lipschitz readouts convert parameter gaps into output deviations. Quantum amplitude estimation adds additive risk certificates with query complexity $\widetilde{\mathcal{O}}(1/\varepsilon\log(1/\delta))$ [61]. Hybrid QEKLR decouples quantum features from convex readouts, enabling post-hoc checks via margins and ROC curves [62].
Differential privacy (DP) complements these by bounding leakage under post-processing. For PQCs, clip parameter-shift gradients, add Gaussian noise at the optimizer or aggregator, and track composition with a moments accountant. Kernel pipelines privatize rows and aggregated statistics. At the client level, clipping plus secure aggregation provides federated privacy [59, 60]. Explainability and QAE-derived uncertainty support DP audits, while Shapley analyses confirm removal at block and gate granularity [61, 58, 63].
@@ 3.3 Mechanisms: Reset, Fisher Guidance, and Gradient Reversal
Mechanisms enact the edit. Reset + partial retraining reinitializes high-influence subspaces and fine-tunes on $D_{s}$. QcNet-like hybrids reset quantum front-ends while retaining classical back-ends [64]; regression-style QCL models exploit few-shot re-optimization [65]. These strategies are shallow and NISQ-friendly.
@@
reduces global distortion while achieving local removal. Empirically, Fisher-aware hybrids outperform random resets and pair well with boosting-style reweighting (AdaBoost.Q) [66].
Gradient reversal and inversion target correlations directly. Parameterized quantum combs approximate inverse processes and minimize Bures divergence to unwind unwanted effects [67]. Ensembles combine reversal with Fisher rebalancing to preserve accuracy under component erasure [66]. Libraries such as sQUlearn package QFIM access, reversal hooks, and reset utilities for practical deployments [68].

### v2

@@ 1 Introduction
Quantum computing and machine intelligence are converging to redefine how data-driven systems learn, adapt, and comply with privacy laws. As quantum machine learning (QML) transitions from small-scale prototypes to distributed deployments, the notion of the right to erasure acquires a new dimension. Under frameworks such as the GDPR, erasure must extend beyond model parameters to include quantum states, channels, and measurements [cao2015towards, bourtoule2021machine]. In this setting, unlearning is not a simple rollback, it is a physically valid transformation that decouples private signals from global quantum behaviour. Quantum mechanics offers natural privacy safeguards through no-cloning, measurement disturbance, and decoherence, yet these same principles make verification, audit, and decentralized enforcement challenging at scale.
@@
Federated learning (FL) decentralizes training but still exposes gradients and aggregates to inversion risks. Quantum Federated Learning (QFL) pushes privacy deeper into the computational stack by encoding updates as quantum states and coordinating learning through variational circuits [qiao2025transitioning]. When combined with quantum differential privacy (QDP), QFL can achieve rigorous confidentiality with minimal utility loss. Rofougaran et al. [rofougaran2024federateddp] demonstrate a QFL–QDP pipeline achieving $98\%$ accuracy under $\epsilon<1.3$ on NISQ hardware, showing that privacy and performance can coexist. This privacy posture now extends to the transport layer. Quantum cloud infrastructures and 6G space–ground networks rely on quantum key distribution (QKD) and homomorphic encryption to guarantee that deleted states remain computationally and physically unrecoverable. Zeng et al. introduce quantum circuit random reconstruction homomorphic encryption (QCRRA), enabling encrypted circuit execution even on untrusted servers [zeng2026quantumcloud]. Similarly, hybrid QFL combined with QKD improves both anomaly detection and key generation rates, indicating that privacy-preserving protocols can enhance, not hinder, performance [mujlid2025quantumiot].
@@
Quantum deployments also expand the threat model. Adversaries may conduct membership inference, inversion, or poisoning attacks on parametrized quantum circuits, exploiting universal perturbations that transfer across architectures [qiu2023uap]. A defensible system must reason about privacy at the level of quantum channels. A natural analogue of $(\varepsilon,\delta)$-differential privacy requires that, for neighbouring datasets encoded as states and for any POVM, the resulting output distributions differ by at most $(\varepsilon,\delta)$ after the full CPTP pipeline. Composition across rounds and clients then follows from channel algebra. In practice, calibrated dephasing or depolarizing noise acts as privacy randomization, while error mitigation restores performance. In federated setups, cryptographic transport is indispensable. Measurement-device-independent QKD provides composable keys for authenticated encryption and Pauli one-time pads during aggregation, as illustrated in smart-grid benchmarks [ren2024qqfl]. Privacy must also persist through time. Quantum continual learning reduces catastrophic forgetting and constrains leakage from past tasks by allocating disjoint PQC subspaces or projecting gradients to avoid interference [xu2024dynamic, situ2023qcl]. Robustness-aware training, randomized encodings, and adversarial PQC optimization further limit attack transferability and restrict what can be inferred from exchanged statistics.
@@
Contributions. (1) A quantum information formalization of unlearning that replaces destructive erasure with CPTP-driven redistribution, anchored in the data-processing inequality and thermodynamic irreversibility. (2) A taxonomy and evaluation framework covering sample-, class-, and client-level scopes; empirical, certified, and DP guarantees; and mechanisms including reset, Fisher-weighted updates, and gradient reversal across QML and QFU deployments. (3) Federated QMU protocols that combine gradient hiding, differential privacy accounting, and cryptographic transport for secure orchestration [li2024privacy, bhatia2024irps, mashetty2025pqfl, ballester2025qflsurvey]. (4) Threat-informed design principles addressing transfer attacks and catastrophic forgetting through continual-learning and robustness-aware PQC training [qiu2023uap, xu2024dynamic, situ2023qcl]. (5) Empirical evidence that privacy and utility can coexist in practice, including hybrid recurrent and time-series models where variational quantum embeddings improve forecasting with compact architectures [yu2023qlstm].
@@ 2.1 From Classical to Quantum Unlearning
Machine unlearning (MU) aims to remove the influence of selected training data without retraining the entire model. The idea aligns with legal and ethical frameworks such as the European Union’s Right to Erasure in the GDPR [cao2015towards, bourtoule2021machine], which demand that data and its effects be verifiably erased. Early approaches used practical engineering strategies, including checkpoint rollback and incremental updates. Influence functions introduced a more principled approach by estimating each sample’s contribution to the optimum through $\mathbf{H}^{-1}\nabla_{\theta}L(\theta^{*},x_{i})$, with $\mathbf{H}$ as the empirical Hessian [koh2017understanding]. The SISA framework (Sharded, Isolated, Sliced Aggregation) further improved auditability and deletion latency through modular retraining [bourtoule2021machine].
Three common goals unify this body of work. Efficiency demands minimal computational cost and wall-clock time. Completeness ensures both data and its downstream influence are removed. Verifiability requires clear, measurable evidence of forgetting. Certified low-rank corrections enable exact updates in convex models [guo2020certified], while approximate regularization and Bayesian re-weighting extend unlearning to deep, non-convex models by transitioning from $p(\theta|D)$ to $p(\theta|D\setminus D_{r})$ [nguyen2022bayesian].
Federated unlearning applies similar ideas to distributed learning. Each client locally compensates for removed data by adjusting its gradient contribution, e.g., $\theta^{\prime}=\theta-\alpha\nabla_{\theta}L_{i}$, often using stored or compressed updates [liu2021federated]. Hybrid quantum–classical systems now adopt comparable coordination strategies. Quantum feature extractors integrate with classical aggregators [bhatia2023federated], ring topologies support bandwidth-efficient communication [wang2024quantum], and dynamic aggregation improves robustness under non-IID healthcare data [qu2025daqfl]. Information-theoretically, ideal unlearning reduces mutual information between the removed data and the model, $I(D_{r};M)$, while retaining $I(D_{s};M)$ for the preserved set [ginart2019making]. Metrics such as forgetting accuracy, retention, and retraining cost quantify progress toward this goal.
@@ 2.2 Quantum Learning Dynamics
which allows exact differentiation under real hardware noise [schuld2019evaluating].
Model expressivity depends on data encoding and circuit depth. Re-uploading strategies expand hypothesis space by alternating data injection with trainable layers [fan2022compact], while differentiable quantum programming provides compiler-level support for gradient computation across control structures [zhu2020pldi]. Training remains challenging due to barren plateaus, where gradient variance decays exponentially with circuit depth [mcclean2018barren]. Geometry-aware optimization addresses this by using the quantum Fisher information (QFI) to rescale gradients, $\theta\leftarrow\theta-\eta F(\theta)^{-1}\nabla_{\theta}\mathcal{L}$, thereby improving convergence and shot efficiency on NISQ devices [stokes2020quantum].
Distributed learning frameworks extend these foundations. In quantum federated learning (QFL), multiple clients train local PQCs or hybrid models and share updates under privacy-preserving protocols such as gradient hiding, quantum differential privacy, or entanglement-based coordination [li2024privacy, bhatia2024irps, mashetty2025pqfl]. Comprehensive surveys document the architectural and privacy challenges that shape scalable QFL systems [ballester2025qflsurvey]. Collectively, exact gradient evaluation, QFI-guided optimization, and privacy-preserving coordination provide the algorithmic substrate on which QMU is built. They allow a quantum model to not only learn from data but also to forget it through controlled, measurable influence reduction.
@@ 2.3 Information Constraints: No-Cloning and No-Deletion
Logical erasure also has energetic cost. Landauer’s principle states that each bit of erased information dissipates heat [landauer1961irreversibility, bennett2003notes]. In QMU, forgetting increases entropy in the environment rather than annihilating stored information. For any completely positive trace-preserving (CPTP) map $\mathcal{E}$ and states $\rho,\sigma$, the data-processing inequality,
@@
ensures that distinguishability monotonically decreases under physical evolution [lindblad1975entropy, uhlmann1977relative]. Trace distance and fidelity exhibit the same contractive behavior, forming the mathematical foundation for certified quantum forgetting.
@@
These physical limits redefine the purpose of unlearning. QMU cannot erase information; it instead makes the information operationally inaccessible. Practical pipelines combine three elements. First, CPTP channels provably contract trace distance or increase fidelity relative to the retrained counterfactual. Second, QFI-guided optimization reduces parameter sensitivity to individual samples. Third, federated and secure protocols prevent reconstruction through gradient hiding or differential privacy [joshi2023blind, shenoy2020measurement, chen2024robust, li2024privacy, ren2024tase]. Adversarial analyses confirm that measurable contraction, not ideal erasure, defines robust unlearning under realistic noise [kundu2025glsvlsi, ergu2025qpoison].
@@ 3.1 Scope: Sample, Class, and Client Levels
Scope specifies the unit of forgetting. At the sample level, the target set $\mathcal{S}=\{x_{j}\}$ contains individual records whose influence should be removed while preserving utility on retained data. In kernel QML, this often reduces to low-rank modifications of the Gram matrix $K_{jk}=\kappa_{\phi}(x_{j},x_{k})$ and efficient Sherman–Morrison–Woodbury updates; shallow, aligned kernels tend to dominate generalization and admit fast row/column re-estimation [wang2025qkar]. In variational models, edits localize to layers of $U(\theta)=\prod_{\ell}U_{\ell}(\theta_{\ell})$ guided by sensitivity; a first-order approximation
@@
captures the local effect of $\mathcal{S}$, with $H^{-1}$ replaced in practice by damped or block-diagonal QFI preconditioners for stability on NISQ hardware [daka2025nv].
At the class level, we remove geometry associated with one or more labels and retain the rest. Mixed-state generative classifiers enable component-level deletion without retraining entire models [useche2025qgc]. Search over circuit families with tunable expressivity supports targeted suppression of class-specific features [beaudoin2025qfusion]. Compact QVC ensembles further allow pruning of subcircuits and light retraining of combiner layers for sequential tasks [moll2025cyber].
At the client level, relevant to QFU, forgetting combines parameter edits with state-theoretic maps. A local CPTP channel acts on the client register, followed by partial retraining of the global model; distributed QLSTM and kernelized pipelines implement such workflows with secure coordination [chen2025dqlstm, hsu2025qklstm, mantha2025pilot]. Applications in security and materials show that lightweight hybrids with small latent spaces permit rapid reinitialization and refitting [sridevi2025hqc, frehner2025qae-ts, kang2025mtv]. Choosing scope is thus a design decision: sample-level for precise removal [wang2025qkar], class-level for generative or structural shifts [useche2025qgc], and client-level when privacy or provenance dominates [chen2025dqlstm, mantha2025pilot]. Auditability and secure aggregation constraints should inform this choice [moreau2025qai, aizpurua2025vqa].
@@ 3.2 Guarantees: Empirical, Certified, and Differential Privacy
Guarantees make forgetting credible. Empirical guarantees rely on reproducible trends across seeds, backends, and datasets. QSVM and kernel hybrids show stable accuracy on compressed medical datasets such as CompressedMediQ and QProteoML [chen2025compressedmediq, priyadharshini2025qproteoml], and robustness persists under decoherence and backend variability [mahdian2025qsvment, khalil2025pegasosqsvm]. These studies validate observed forgetting behavior, although they do not bound worst cases.
@@
strengthened by precompression [chen2025compressedmediq]. For variational models,
@@
and Lipschitz readouts convert parameter gaps into output deviations. Quantum amplitude estimation adds additive risk certificates with query complexity $\widetilde{\mathcal{O}}(1/\varepsilon\log(1/\delta))$ [delejarza2025qint]. Hybrid QEKLR decouples quantum features from convex readouts, enabling post-hoc checks via margins and ROC curves [misra2025qeklr].
Differential privacy (DP) complements these by bounding leakage under post-processing. For PQCs, clip parameter-shift gradients, add Gaussian noise at the optimizer or aggregator, and track composition with a moments accountant. Kernel pipelines privatize rows and aggregated statistics. At the client level, clipping plus secure aggregation provides federated privacy [mahdian2025qsvment, khalil2025pegasosqsvm]. Explainability and QAE-derived uncertainty support DP audits, while Shapley analyses confirm removal at block and gate granularity [delejarza2025qint, priyadharshini2025qproteoml, heese2025shapley].
@@ 3.3 Mechanisms: Reset, Fisher Guidance, and Gradient Reversal
Mechanisms enact the edit. Reset + partial retraining reinitializes high-influence subspaces and fine-tunes on $D_{s}$. QcNet-like hybrids reset quantum front-ends while retaining classical back-ends [gunasekar2025qcnet]; regression-style QCL models exploit few-shot re-optimization [mohamed2025qcl]. These strategies are shallow and NISQ-friendly.
@@
reduces global distortion while achieving local removal. Empirically, Fisher-aware hybrids outperform random resets and pair well with boosting-style reweighting (AdaBoost.Q) [chen2025adaboostq].
Gradient reversal and inversion target correlations directly. Parameterized quantum combs approximate inverse processes and minimize Bures divergence to unwind unwanted effects [mo2025pqcomb]. Ensembles combine reversal with Fisher rebalancing to preserve accuracy under component erasure [chen2025adaboostq]. Libraries such as sQUlearn package QFIM access, reversal hooks, and reset utilities for practical deployments [kreplin2025squlearn].